REINHARDT DOMAINS WITH

NON-COMPACT AUTOMORPHISM GROUPS

Siqi Fu Alexander V. Isaev Steven G. Krantz

Abstract. We give an explicit description of smoothly bounded Reinhardt domains with noncompact automorphism groups. In particular, this description confirms a special case of a conjecture of Greene/Krantz.

Typeset by AMS-TEX